10 years from now Jane will be 3 times older than she was 4 years ago. How old is Jane
Answer by macston(5194)

J=Jane's age
J+10yrs=3(J-4yrs)
J+10yrs=3J-12yrs Add12yrs-J to each side.
22yrs=2J Divide each side by 2.
11yrs=J ANSWER: Jane is 11 years old.